OpenStudy (anonymous):

D = M/V solve for M. D = 5.1 and V = 0.3

OpenStudy (anonymous):

So 5.1 =M/ 0.3

OpenStudy (anonymous):

Rewriting Equations

OpenStudy (anonymous):

m=5.1*0.3 or m=d*v

OpenStudy (anonymous):

never mind people